OpenStudy (xmoses1):

The slope of the line whose equation is 3y + 2x = 1 is -2/3 -3/2 1/3

jigglypuff314 (jigglypuff314):

you can try putting it into slope intercept form: y=mx+b where m=slope to do that for 3y + 2x = 1 subtract 2x from both sides then divide both sides by 3

OpenStudy (xmoses1):

oh ok then the answer is 1/3

jigglypuff314 (jigglypuff314):

not quite :) the slope intercept form of 3y + 2x = 1 would be y = -2/3 x + 1/3 ^ this part is the slope :)

OpenStudy (xmoses1):

ohhh... oops Xp

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (xmoses1):

thank you @jigglypuff314 :)